A is 20years older than b.5years ago,a was 3 times as old as b.find their present ages.

Answer by stanbon(75887)

A is 20 years older than b. 5years ago,a was 3 times as old as b.
find their present ages.
--------
Equations:
a = b + 20
a-5 = 3(b-5)
----
Substitute for "a" and solve for "b":
b+20-5 = 3(b-5)
---
b+15 = 3b-15
2b = 30
b = 15 yrs (B's age now)
--------
Solve for "a" using a = b + 20
a = 15+20
a = 35 yrs (A's age now)
==================
Cheers,
Stan H.
====================